# IDEATOR: Jailbreaking and Benchmarking Large Vision-Language Models Using Themselves

### v3

@@ Abstract
As large Vision-Language Models (VLMs) gain prominence, ensuring their safe deployment has become critical. Recent studies have explored VLM robustness against jailbreak attacks—techniques that exploit model vulnerabilities to elicit harmful outputs. However, the limited availability of diverse multimodal data has constrained current approaches to rely heavily on adversarial or manually crafted images derived from harmful text datasets, which often lack effectiveness and diversity across different contexts. In this paper, we propose IDEATOR, a novel jailbreak method that autonomously generates malicious image-text pairs for black-box jailbreak attacks. IDEATOR is grounded in the insight that VLMs themselves could serve as powerful red team models for generating multimodal jailbreak prompts. Specifically, IDEATOR leverages a VLM to create targeted jailbreak texts and pairs them with jailbreak images generated by a state-of-the-art diffusion model. Extensive experiments demonstrate IDEATOR’s high effectiveness and transferability, achieving a 94% attack success rate (ASR) in jailbreaking MiniGPT-4 with an average of only 5.34 queries, and high ASRs of 82%, 88%, and 75% when transferred to LLaVA, InstructBLIP, and Chameleon, respectively. Building on IDEATOR’s strong transferability and automated process, we introduce the VLBreakBench, a safety benchmark comprising 3,654 multimodal jailbreak samples. Our benchmark results on 11 recently released VLMs reveal significant gaps in safety alignment. For instance, our challenge set achieves ASRs of 46.31% on GPT-4o and 19.65% on Claude-3.5-Sonnet, underscoring the urgent need for stronger defenses. Disclaimer: This paper contains content that may be disturbing or offensive.
@@ 1 Introduction
In our setup, the attacker VLM acts as an “ideator” that simulates an adversary interacting with the target VLM. The attacker VLM iteratively refines its strategy based on the target’s previous responses, while the target VLM processes only the current input without access to historical conversations. IDEATOR also employs concurrent attack streams, exploring multiple jailbreak strategies simultaneously, enabling a comprehensive examination of VLM vulnerabilities. Furthermore, IDEATOR’s ability to autonomously generate diverse and contextually rich attack samples supports large-scale, cross-model evaluations, making it a crucial tool for assessing VLM safety. Using IDEATOR, we introduce VLBreakBench, a safety benchmark consisting of 3,654 multimodal jailbreak samples, and perform evaluations on 11 recently released VLMs.
@@
  • Using IDEATOR, we construct a multimodal safety benchmark named VLBreakBench for VLMs, which consists of 3,654 multimodal jailbreak samples. Evaluations on 11 state-of-the-art VLMs reveal significant gaps in current safety mechanisms, underscoring the need for stronger defenses.
@@ 2.1 Large Vision-Language Models (VLMs)
Large VLMs extend traditional Large Language Models (LLMs) by integrating visual and textual modalities. Typically, VLMs combine a pre-trained LLM with an image encoder, mapping visual features to the LLM’s token space via an alignment module. For example, MiniGPT-4 [40] aligns a frozen visual encoder [12] with a frozen LLM [9] using a single projection layer, while InstructBLIP [10] introduces an instruction-aware Query Transformer to extract task-relevant features. LLaVA [20] connects a vision encoder [27] with an LLM [33], leveraging GPT-4-generated multimodal data [1] for instruction tuning. Despite their advanced capabilities, the integration of visual modalities introduces new vulnerabilities [29, 19, 18, 38], highlighting the need for robust alignment strategies.
@@ 2.2 Jailbreak Attacks on VLMs
Recent studies have explored various VLM jailbreak strategies. Greshake et al. [16] injected deceptive text into images, while Gong et al. [15] proposed FigStep, converting harmful text into images to bypass safeguards. Liu et al. [21] showed VLMs can be compromised by query-relevant images and introduced MM-SafetyBench for robustness evaluation. Other works [3, 4, 6] focus on adversarial optimization approaches, which optimize adversarial images to increase the likelihood of attacker-chosen text outputs. Similarly, the Visual Adversarial Jailbreak Method (VAJM) [26] used a single adversarial image to universally jailbreak aligned VLMs, generating harmful content beyond the initial optimization scope. Recently, Wang et al. [35] introduced a dual-modality attack that simultaneously generates adversarial image prefixes and text suffixes, enabling more sophisticated and effective jailbreaks. Niu et al. [25] extended this approach by transforming adversarial images into text suffixes for LLMs.
@@ 3.2.3 Breadth-Depth Exploration
8:              $\mathcal{O}_{\text{json}}^{(b,d)}&=\mathcal{M}_{\mathcal{A}}(I_{b,d-1},\mathcal{R}_{b,d-1})\\ &=\{\mathcal{A}_{b,d},P_{t}^{(b,d)},P_{i}^{(b,d)}\}\end{aligned}$
@@ 3.2.4 Attacker Model Selection
VLBreakBench Construction To systematically evaluate the safety of both open-source and commercial VLMs against multimodal jailbreak attacks, we introduce VLBreakBench, a benchmark constructed using diverse multimodal jailbreak prompts generated by IDEATOR. To enhance the stealth and sophistication of the attack samples, we further adapt IDEATOR to Google’s Gemini [31] with its safety settings disabled. Leveraging Gemini as a stronger base model significantly improves jailbreak effectiveness, particularly against more secure commercial VLMs, thereby creating a more challenging evaluation set for VLBreakBench. For high-quality image generation, we employ Stable Diffusion 3.5 Large [28]. VLBreakBench enables a rigorous and adversarial assessment of VLM vulnerabilities, providing a comprehensive framework for evaluating model robustness.
@@ 4 IDEATOR Evaluation Experiments
In this section, we first describe the experimental setup and then present the evaluation results of IDEATOR, focusing on its attack effectiveness and transferability to other VLMs. The detailed construction of VLBreakBench and the benchmarking experiments are deferred to Section 5.
@@ 4.4 Visualization and Ablation
$N_{\text{d}}$ $N_{\text{b}}$ 1 3 5 7 Attack Type ASR (%) Avg. #Queries
@@
5 VLBreakBench
Given the rich semantic content, high transferability, and diversity of the generated jailbreak samples, IDEATOR serves as an ideal tool for constructing a VLM safety benchmark. While existing benchmarks [37, 21] primarily focus on detecting explicit harmful content, they lack the capability to evaluate VLM robustness against sophisticated multimodal jailbreak attacks. To address this gap, we introduce VLBreakBench, a novel benchmark specifically designed to assess VLMs in realistic adversarial scenarios. It provides a comprehensive and practical evaluation of VLM vulnerabilities, enabling a deeper understanding of their safety limitations in real-world applications.
@@ 5.1 Benchmark Overview
VLBreakBench is structured into two evaluation tiers: a base set and a challenge set, designed to assess VLMs at distinct difficulty levels. The dataset spans 12 safety topics and 46 subcategories, comprising 916 harmful queries. For each query, we generate one jailbreak text-image pair for the base set and three for the challenge set, resulting in a comprehensive collection of 3,654 jailbreak samples. This hierarchical design ensures a rigorous evaluation of VLM robustness across varying adversarial scenarios.
Figure 6: Safety taxonomy of VLBreakBench.
Safety Risk Taxonomy  To construct a comprehensive safety risk taxonomy for VLBreakBench, we collaborate with experts from the humanities and social sciences to extend existing taxonomies [21], ensuring coverage of both technical vulnerabilities and societal impacts. Figure 6 presents our taxonomy, while Appendix D provides detailed secondary classifications, a statistical overview, and representative attack examples from the challenge set.
Table 4: Safety evaluation of 11 VLMs on the challenge set of VLBreakBench, measured by ASR (%) across 12 safety topics. Safety topics and certain model names are abbreviated for brevity. “Avg.” denotes the average ASR across all topics.
@@ 5.3 Benchmarking Results
We evaluate 11 state-of-the-art VLMs, including both open-source and commercial models: MiniGPT-v2 (Llama-2-Chat-7B) [8], LLaVA-OneVision (7B) [17], Qwen2-VL (7B) [34], Llama-3.2-11B/90B-Vision-Instruct [13], Gemini-1.5-Pro [32], Gemini-2.0-Flash/Flash-Thinking [11], GPT-4o Mini [1], GPT-4o [1], and Claude-3.5-Sonnet [2]. All commercial models use their latest versions as of February 2025. The ASR evaluation is automated using Gemini-2.0-Flash-Thinking. Table 4 summarizes the challenge set results across the 11 VLMs, with base set results provided in Appendix E.
@@
Limitations While our IDEATOR proves effective in automating jailbreaks using accessible VLMs and diffusion models, its utility is constrained by the trade-off between the weak alignment and strong capabilities of attacker models. Additionally, although VLBreakBench serves as a useful benchmark for multimodal safety, its current scale is relatively small, necessitating further computational resources and automated selection methods to expand its scope.
@@ 6 Conclusion
In this paper, we proposed a novel black-box jailbreak method IDEATOR to uncover safety vulnerabilities in large VLMs. By utilizing VLMs as red team models, IDEATOR autonomously generates adversarial image-text pairs, offering a scalable framework for safety evaluation. Experiments demonstrate IDEATOR’s effectiveness and transferability, achieving a 94% success rate in jailbreaking MiniGPT-4 with an average of only 5.34 queries, and high transfer success rates of 82%, 88%, and 75% on LLaVA, InstructBLIP, and Meta’s Chameleon, respectively. Leveraging IDEATOR’s capabilities, we constructed a multimodal jailbreak benchmark VLBreakBench to evaluate VLMs against diverse adversarial scenarios, differentiating itself from exisint non-adversarial multimodal safety benchmarks. Benchmarking 11 state-of-the-art VLMs on 3,654 multimodal jailbreak samples reveals significant safety gaps, with GPT-4o and Claude-3.5-Sonnet achieving attack success rates of 46.31% and 19.65%, respectively. Future work will involve scaling up VLBreakBench and exploring defense methods utilizing IDEATOR.
@@ C Additional Visualization Results
D VLBreakBench Data Statistics
Table 6 presents a statistical summary of the base set and challenge set in VLBreakBench, organized by safety-critical categories and their subcategories. The dataset spans a wide range of safety risks, including economic harm, ethical issues, hate speech, cyberattacks, privacy violations, and violent behavior, offering a comprehensive evaluation of VLM robustness against multimodal threats. Figure 12 showcases examples from the challenge set, which feature more subtle and sophisticated attacks designed to test the limits of current defenses.
@@ D VLBreakBench Data Statistics
Figure 12: Examples from the challenge set in VLBreakBench. These examples showcase the types of complex scenarios used to test the robustness of VLMs.
E Benchmarking Results on the Base Set
Table 7 presents the safety evaluation results of 11 VLMs on the base set of VLBreakBench. The base set assesses fundamental vulnerabilities in VLMs. Among open-source models, Qwen2-VL exhibits the highest vulnerability with an average ASR of 35.04%. In contrast, Llama-3.2-90B-Vision demonstrates the strongest robustness among open-source models, with an average ASR of 7.97%. For commercial models, Gemini-2.0-Flash is the most vulnerable, with an ASR of 53.38%, performing worse than some open-source alternatives. Claude-3.5-Sonnet remains the most robust overall, with an ASR as low as 1.09%, significantly outperforming all other models.
Table 7: Safety evaluation of 11 VLMs on the base set of VLBreakBench, measured by ASR across 12 safety topics. Safety topics and certain model names are abbreviated for brevity. “Avg.” denotes the average ASR across all topics.

### v4

@@ Abstract
As large Vision-Language Models (VLMs) gain prominence, ensuring their safe deployment has become critical. Recent studies have explored VLM robustness against jailbreak attacks—techniques that exploit model vulnerabilities to elicit harmful outputs. However, the limited availability of diverse multimodal data has constrained current approaches to rely heavily on adversarial or manually crafted images derived from harmful text datasets, which often lack effectiveness and diversity across different contexts. In this paper, we propose IDEATOR, a novel jailbreak method that autonomously generates malicious image-text pairs for black-box jailbreak attacks. IDEATOR is grounded in the insight that VLMs themselves could serve as powerful red team models for generating multimodal jailbreak prompts. Specifically, IDEATOR leverages a VLM to create targeted jailbreak texts and pairs them with jailbreak images generated by a state-of-the-art diffusion model. Extensive experiments demonstrate IDEATOR’s high effectiveness and transferability, achieving a 94% attack success rate (ASR) in jailbreaking MiniGPT-4 with an average of only 5.34 queries, and high ASRs of 82%, 88%, and 75% when transferred to LLaVA, InstructBLIP, and Chameleon, respectively. Building on IDEATOR’s strong transferability and automated process, we introduce the VLJailbreakBench, a safety benchmark comprising 3,654 multimodal jailbreak samples. Our benchmark results on 11 recently released VLMs reveal significant gaps in safety alignment. For instance, our challenge set achieves ASRs of 46.31% on GPT-4o and 19.65% on Claude-3.5-Sonnet, underscoring the urgent need for stronger defenses. Disclaimer: This paper contains content that may be disturbing or offensive.
@@ 1 Introduction
In our setup, the attacker VLM acts as an “ideator” that simulates an adversary interacting with the target VLM. The attacker VLM iteratively refines its strategy based on the target’s previous responses, while the target VLM processes only the current input without access to historical conversations. IDEATOR also employs concurrent attack streams, exploring multiple jailbreak strategies simultaneously, enabling a comprehensive examination of VLM vulnerabilities. Furthermore, IDEATOR’s ability to autonomously generate diverse and contextually rich attack samples supports large-scale, cross-model evaluations, making it a crucial tool for assessing VLM safety. Using IDEATOR, we introduce VLJailbreakBench, a safety benchmark consisting of 3,654 multimodal jailbreak samples, and perform evaluations on 11 recently released VLMs.
@@
  • Using IDEATOR, we construct a multimodal safety benchmark named VLJailbreakBench for VLMs, which consists of 3,654 multimodal jailbreak samples. Evaluations on 11 state-of-the-art VLMs reveal significant gaps in current safety mechanisms, underscoring the need for stronger defenses.
@@ 2.1 Large Vision-Language Models (VLMs)
Large VLMs extend traditional Large Language Models (LLMs) by integrating visual and textual modalities. Typically, VLMs combine a pre-trained LLM with an image encoder, mapping visual features to the LLM’s token space via an alignment module. For example, MiniGPT-4 [44] aligns a frozen visual encoder [12] with a frozen LLM [9] using a single projection layer, while InstructBLIP [10] introduces an instruction-aware Query Transformer to extract task-relevant features. LLaVA [23] connects a vision encoder [31] with an LLM [37], leveraging GPT-4-generated multimodal data [1] for instruction tuning. Despite their advanced capabilities, the integration of visual modalities introduces new vulnerabilities [33, 22, 21, 42, 27], highlighting the need for robust alignment strategies.
@@ 2.2 Jailbreak Attacks on VLMs
Recent studies have explored various VLM jailbreak strategies. Greshake et al. [19] injected deceptive text into images, while Gong et al. [18] proposed FigStep, converting harmful text into images to bypass safeguards. Liu et al. [24] showed VLMs can be compromised by query-relevant images and introduced MM-SafetyBench for robustness evaluation. Other works [3, 4, 6, 14] explore adversarial optimization techniques that generate adversarial images by either maximizing the likelihood of attacker-specified outputs or aligning visual embeddings with those of harmful content. Similarly, the Visual Adversarial Jailbreak Method (VAJM) [30] used a single adversarial image to universally jailbreak aligned VLMs, generating harmful content beyond the initial optimization scope. Recently, Wang et al. [39] introduced a dual-modality attack that simultaneously generates adversarial image prefixes and text suffixes, enabling more sophisticated and effective jailbreaks. Niu et al. [29] extended this approach by transforming adversarial images into text suffixes for LLMs.
@@ 3.2.3 Breadth-Depth Exploration
8:     $\begin{aligned} \mathcal{O}_{\text{json}}^{(b,d)}&=\mathcal{M}_{\mathcal{A}}(I_{b,d-1},\mathcal{R}_{b,d-1})\\ &=\{\mathcal{A}_{b,d},P_{t}^{(b,d)},P_{i}^{(b,d)}\}\end{aligned}$
@@ 3.2.4 Attacker Model Selection
VLJailbreakBench Construction To systematically evaluate the safety of both open-source and commercial VLMs against multimodal jailbreak attacks, we introduce VLJailbreakBench, a benchmark constructed using diverse multimodal jailbreak prompts generated by IDEATOR. To enhance the stealth and sophistication of the attack samples, we further adapt IDEATOR to Google’s Gemini [35] with its safety settings disabled. Leveraging Gemini as a stronger base model significantly improves jailbreak effectiveness, particularly against more secure commercial VLMs, thereby creating a more challenging evaluation set for VLJailbreakBench. For high-quality image generation, we employ Stable Diffusion 3.5 Large [32]. VLJailbreakBench enables a rigorous and adversarial assessment of VLM vulnerabilities, providing a comprehensive framework for evaluating model robustness.
@@ 4 IDEATOR Evaluation Experiments
In this section, we first describe the experimental setup and then present the evaluation results of IDEATOR, focusing on its attack effectiveness and transferability to other VLMs. The detailed construction of VLJailbreakBench and the benchmarking experiments are deferred to Section 5.
@@ 4.4 Visualization and Ablation
$N_{\text{d}}$ $N_{\text{b}}$ 1 3 5 7
Attack
Type
ASR
Avg.
#Queries
@@
5 VLJailbreakBench
Given the rich semantic content, high transferability, and diversity of the generated jailbreak samples, IDEATOR serves as an ideal tool for constructing a VLM safety benchmark. Existing benchmarks predominantly target either explicit harmful content detection [41, 24] or focus on text-based transfer attacks [15], leaving a critical gap in assessing the robustness of VLMs against sophisticated multimodal jailbreak threats. To address this gap, we introduce VLJailbreakBench, a novel benchmark specifically designed to assess VLMs in realistic adversarial scenarios. It provides a comprehensive and practical evaluation of VLM vulnerabilities, enabling a deeper understanding of their safety limitations in real-world applications.
@@ 5.1 Benchmark Overview
VLJailbreakBench is structured into two evaluation tiers: a base set and a challenge set, designed to assess VLMs at distinct difficulty levels. The dataset spans 12 safety topics and 46 subcategories, comprising 916 harmful queries. For each query, we generate one jailbreak text-image pair for the base set and three for the challenge set, resulting in a comprehensive collection of 3,654 jailbreak samples. This hierarchical design ensures a rigorous evaluation of VLM robustness across varying adversarial scenarios.
Figure 6: Safety taxonomy of VLJailbreakBench.
Safety Risk Taxonomy  To construct a comprehensive safety risk taxonomy for VLJailbreakBench, we collaborate with experts from the humanities and social sciences to extend existing taxonomies [24], ensuring coverage of both technical vulnerabilities and societal impacts. Figure 6 presents our taxonomy, while Appendix E provides detailed secondary classifications, a statistical overview, and representative attack examples from the challenge set.
Table 4: Safety evaluation of 11 VLMs on the challenge set of VLJailbreakBench, measured by ASR (%) across 12 safety topics. Safety topics and certain model names are abbreviated for brevity. “Avg.” denotes the average ASR across all topics.
@@ 5.3 Benchmarking Results
We evaluate 11 state-of-the-art VLMs, including both open-source and commercial models: MiniGPT-v2 (Llama-2-Chat-7B) [8], LLaVA-OneVision (7B) [20], Qwen2-VL (7B) [38], Llama-3.2-11B/90B-Vision-Instruct [13], Gemini-1.5-Pro [36], Gemini-2.0-Flash/Flash-Thinking [11], GPT-4o Mini [1], GPT-4o [1], and Claude-3.5-Sonnet [2]. All commercial models use their latest versions as of February 2025. The ASR evaluation is automated using Gemini-2.0-Flash-Thinking. Table 4 summarizes the challenge set results across the 11 VLMs, with base set results provided in Appendix F.
@@
Limitations While IDEATOR proves effective in automating jailbreaks using accessible VLMs and diffusion models, its utility is constrained by the trade-off between the weak alignment and strong capabilities of attacker models. Additionally, although VLJailbreakBench serves as a useful benchmark for multimodal safety, its current scale is relatively small, necessitating further computational resources and automated selection methods to expand its scope.
@@ 6 Conclusion
In this paper, we propose IDEATOR, a novel black-box jailbreak method for uncovering safety vulnerabilities in VLMs. By utilizing VLMs as red team models, IDEATOR autonomously generates adversarial image-text pairs, offering a scalable framework for safety evaluation. Experiments demonstrate IDEATOR’s effectiveness and transferability, achieving a 94% success rate in jailbreaking MiniGPT-4 with an average of only 5.34 queries, and high transfer success rates of 82%, 88%, and 75% on LLaVA, InstructBLIP, and Chameleon, respectively. Building on IDEATOR, we construct VLJailbreakBench to evaluate VLMs against diverse adversarial scenarios, differentiating itself from existing safety benchmarks. Benchmarking 11 state-of-the-art VLMs on 3,654 multimodal jailbreak samples reveals significant safety gaps, with GPT-4o and Claude-3.5-Sonnet achieving attack success rates of 46.31% and 19.65%, respectively. Code and benchmark are available at https://github.com/roywang021/IDEATOR and https://huggingface.co/datasets/wang021/VLBreakBench.
Acknowledgements
This work is in part supported by National Key R&D Program of China (Grant No. 2022ZD0160103) and National Natural Science Foundation of China (Grant No. 62276067).
@@ C Additional Visualization Results
D Evaluation against Defense Mechanisms
To provide additional insights into IDEATOR’s robustness against existing defense mechanisms, we performed preliminary experiments on AdaShield-S [16]. AdaShield-S is a recently proposed defense framework designed specifically to detect and mitigate structure-based jailbreak attacks on VLMs. Table 6 presents the attack success rate (ASR) of IDEATOR and two other state-of-the-art black-box methods, Figstep and MM-SafetyBench, before and after applying AdaShield-S.
As indicated in Table 6, IDEATOR demonstrates strong resilience against AdaShield-S, maintaining high ASRs with minimal performance degradation across all tested victim models. Specifically, IDEATOR’s ASR decreased slightly from 94.0% to 84.0% (–10.0%) on MiniGPT-4, 82.0% to 73.0% (–9.0%) on LLaVA, and 88.0% to 87.0% (–1.0%) on InstructBLIP. In comparison, Figstep and MM-SafetyBench experienced substantially larger reductions in ASR, highlighting IDEATOR’s advantage in generating diverse and subtle jailbreak strategies that effectively evade structure-based detection.
Table 6: ASR before and after applying AdaShield-S [16]. Values in parentheses indicate absolute change in ASR.
MiniGPT-4 LLaVA InstructBLIP
Figstep 69.0 $\rightarrow$ 37.0   (-32.0) 44.0 $\rightarrow$ 6.0      (-38.0) 11.0 $\rightarrow$ 16.0    (+5.0)
$M$-SafetyBench 66.0 $\rightarrow$ 37.0   (-29.0) 46.0 $\rightarrow$ 5.0      (-41.0) 29.0 $\rightarrow$ 27.0    (-2.0)
IDEATOR (Ours) 94.0 $\rightarrow$ 84.0    (-10.0) 82.0 $\rightarrow$ 73.0     (-9.0) 88.0 $\rightarrow$ 87.0    (-1.0)
E VLJailbreakBench Data Statistics
Table 7 presents a statistical summary of the base set and challenge set in VLJailbreakBench, organized by safety-critical categories and their subcategories. The dataset spans a wide range of safety risks, including economic harm, ethical issues, hate speech, cyberattacks, privacy violations, and violent behavior, offering a comprehensive evaluation of VLM robustness against multimodal threats. Figure 12 showcases examples from the challenge set, which feature more subtle and sophisticated attacks designed to test the limits of current defenses.
@@ E VLJailbreakBench Data Statistics
Figure 12: Examples from the challenge set in VLJailbreakBench. These examples showcase the types of complex scenarios used to test the robustness of VLMs.
F Benchmarking Results on the Base Set
Table 8 presents the safety evaluation results of 11 VLMs on the base set of VLJailbreakBench. The base set assesses fundamental vulnerabilities in VLMs. Among open-source models, Qwen2-VL exhibits the highest vulnerability with an average ASR of 35.04%. In contrast, Llama-3.2-90B-Vision demonstrates the strongest robustness among open-source models, with an average ASR of 7.97%. For commercial models, Gemini-2.0-Flash is the most vulnerable, with an ASR of 53.38%, performing worse than some open-source alternatives. Claude-3.5-Sonnet remains the most robust overall, with an ASR as low as 1.09%, significantly outperforming all other models.
Table 8: Safety evaluation of 11 VLMs on the base set of VLJailbreakBench, measured by ASR across 12 safety topics. Safety topics and certain model names are abbreviated for brevity. “Avg.” denotes the average ASR across all topics.